OpenStudy (rainbow_dashie):

A recipe calls for 3 cans of tomato sauce. Fawn wants to make 40% fewer servings than the recipe makes. Which expression represents how many cans of tomato sauce Fawn should use? @iambatman

OpenStudy (rainbow_dashie):

@roz_mar23

OpenStudy (rainbow_dashie):

@Destinymasha

OpenStudy (rainbow_dashie):

@roz_mar23

OpenStudy (rainbow_dashie):

okay can u help

OpenStudy (anonymous):

x/3=.4

OpenStudy (rainbow_dashie):

no its got to be (x times x)

OpenStudy (rainbow_dashie):

\[n \times m\]

OpenStudy (anonymous):

or u can do .4*3=x

OpenStudy (rainbow_dashie):

hold on here are the answer choices.

OpenStudy (rainbow_dashie):

http://prntscr.com/5fbl1e

OpenStudy (anonymous):

Well ur ansewr is 1.2 so check eachone and see what one gets 1.2

OpenStudy (rainbow_dashie):

huh

OpenStudy (anonymous):

The answer is 1.2 cups or 1 2/5... do each equation and find the one that equals 1.2, or 6/5, so A

OpenStudy (rainbow_dashie):

thanks

OpenStudy (rainbow_dashie):

it was wrong